OpenStudy (anonymous):

Evaluate (5 + 2)2 − 40 + 3 = 12??

OpenStudy (anonymous):

i got 49 - 40 + 3= 9 + 13= 12?

OpenStudy (anonymous):

also i'd like to remind you that its (5 + 2)^2 - 40 + 3

OpenStudy (kxttxn):

I think you meant 3, not 13, but you did solve it correctly, as 9+3 does equal 12

OpenStudy (anonymous):

okay ty, once again :)

OpenStudy (kxttxn):

no problem :)